a boy inside a bus is moving 5.0 m/s towards the east. he is walking at 1.0m/s toward the front. what is the velocity of the boy as observed by a passenger sitting inside the bus and another passenger waiting for the bus at a nearby bus stop
1
Expert's answer
2019-08-19T09:52:49-0400

The velocity of the boy as observed by a passenger sitting inside the bus is

"5.0 \frac{m}{s}"

The elocity of the boy as observed by another passenger waiting for the bus at a nearby bus stop:


"5.0+1.0=6.0 \frac{m}{s}"